# Phased Instruction Fine-Tuning for Large Language Models

### v1

@@ Abstract
Instruction Fine-Tuning, a method enhancing pre-trained language models’ capabilities from mere next-word prediction to complex instruction following, often employs a one-off training approach on diverse instruction dataset. However, this method may not effectively enhance models’ adherence to instructions due to the simultaneous handling of varying instruction complexities. To address this, we propose a novel phased instruction fine-tuning (Phased IFT) method, grounded in the hypothesis of progressive alignment, which posits that the transition of a pre-trained language model from simple next-word prediction to sophisticated instruction following is a gradual learning process. Specifically, we obtain the score of difficulty for each instruction via GPT-4, stratify the instruction data into subsets of increasing difficulty, and sequentially uptrain on these subsets using the standard supervised loss. Through extensive experiments on the pre-trained models Llama-2 7B/13B, and Mistral-7B using the 52K Alpaca instruction data, we demonstrate that Phased IFT significantly surpasses traditional one-off instruction fine-tuning (One-off IFT) method in win rate, empirically validating the progressive alignment hypothesis. Our findings suggest that Phased IFT offers a simple yet effective pathway for elevating the instruction-following capabilities of pre-trained language models. Models and datasets from our experiments are freely available at https://github.com/xubuvd/PhasedSFT.
@@ 1 Introduction
Figure 1: Win rate comparison between: (the red line) Phased IFT on difficulty-stratified instruction data and (the yellow line) Phased IFT on randomly stratified instruction data, against (the gray dotted line) a baseline of One-off IFT on the whole instruction dataset.
@@
To address the aforementioned issue, this paper introduces an instruction difficulty scoring mechanism. Specifically, the difficulty of an instruction is defined by the complexity of the instruction itself, the input, and the challenges involved in generating the output. We employ the strongest available models (such as GPT-4 OpenAI (2023)) as a teacher to score the difficulty of each instruction and input, as well as the challenge in generating the output, on a scale from 1 to 5, where higher scores indicate greater difficulty. Further, we plot probability density curve of the instruction difficulty as a heuristic guide Campello et al. (2020), coupled with expert judgment to select thresholds for difficulty scores. This approach enabled us to segment the instruction dataset into multi-stages sub-datasets, forming a sequence with increasing difficulty levels. Building upon this, we further propose the progressive alignment hypothesis: Aligning a pretrained model’s existing ability to predict the next word with the capability to generate content following human intent is a gradual learning process, progressively attaining alignment with human intent.
@@
To assess this method, we engage in extensive experiments using two widely used instruction datasets: Alpaca 52K and its refined variant, Alpaca-clean, in conjunction with three leading open-source pre-trained models: Llama-2 7/13B Touvron et al. (2023) and Mistral-7B Jiang et al. (2023). GPT-4 served as judge to assess the gains of Phased IFT against One-off IFT in win rate. In evaluations across 6 benchmark datasets, substantial enhancements in win rate are recorded for the Phased IFT method. Figure 1 illustrates the win rate as a function of the uptraining stages for phased IFT versus the traditional one-off IFT method. The red line indicates phased IFT on difficulty-staged data, after uptraining in the stage 2, the win rate already surpasses that of the one-off IFT method, and after the stage 3 of uptraining, the win rate significantly increases. The yellow line shows phased IFT on randomly staged data, the win rate after fine-tuning in any stage is lower than that of the traditional one-off IFT method, highlighting the substantial advantage of our proposed phased instruction fine-tuning method. Furthermore, extensive ablation studies validate the hypothesis and the efficacy of the Phased IFT, providing robust empirical support. The main contributions of this paper include:
@@
  • We introduce the progressive alignment hypothesis, suggesting that aligning the generative abilities of pre-trained models with human instructions is a process of incremental training, rather than achievable through one-off fine-tuning. This hypothesis is substantiated through extensive experimental validation.
  • A series of experiments demonstrate that our Phased IFT method is more effective than One-off IFT approach. In terms of win rate across six benchmark, the Phased IFT shows significant improvements.
@@ 2 Related Work
Open-source Instruction-tuned Models Recent studies have shown that fine-tuning Large Language Models (LLMs) with open-source instruction datasets significantly boosts their adherence to human instructions. The 65B parameter LIMA Zhou et al. (2023) model, after fine-tuning with 1,000 human-generated instructions, outperformed GPT-4 in 43% of cases, Bard in 58%, and DaVinci003 in 65%, showcasing its superior instruction-following capability. Similarly, the Alpaca 7B model Taori et al. (2023), fine-tuned on 52K instructions from text-davinci-003, matched its performance, demonstrating the effectiveness of large-scale, diverse dataset fine-tuning. Advancing further, Alpagasus Chen et al. (2024) selects a subset of 9k high-quality instructions from the Alpaca 52K and is fine-tuned on the subset, significantly surpassing the original Alpaca’s Taori et al. (2023) performance across four human-instruction test sets, e.g., Self-Instruct Wang et al. (2023b), WizardLM Xu et al. (2024), Koala Geng et al. (2023) and Vicuna Chiang et al. (2023), highlighting the benefits of precise instruction selection and targeted fine-tuning for enhancing LLM capabilities. Specifically, Alpagasus prompts ChatGPT to score each instruction on a scale of 0 to 5, resulting in a predominant score range of 3 to 4, comprising roughly 78.9% of its total. Then Alpagasus selectively utilizes 9,229 instructions rated above 4.5, establishing a high-quality dataset for fine-tuning.
Open-source Instruction Dataset The Alpaca dataset Taori et al. (2023), designed to improve the instruction-following capabilities of language models and comprising 52K instructions, was generated using OpenAI’s text-davinci-003. This refined the Self-Instruct Wang et al. (2023b) framework, resulting in a dataset with significantly greater diversity than its Self-Instruct counterpart. Alpaca-cleaned yahma (2023) is a cleaned version of the original Alpaca dataset, identified several issues, such as hallucinations, empty outputs and wrong answers, thereby elevating the overall quality of the data. Additionally, LIMA Zhou et al. (2023) released a meticulously curated dataset comprising 1000 instructions, incorporating 750 popular queries and responses from community forums like Stack Exchange and wikiHow, alongside 250 expertly crafted entries. This compilation emphasizes the quality and diversity of instructions.
Instruction Fine-tuning Strategy The prevalent method of instruction fine-tuning, employed in models such as LIMA, Alpaca, Alpagasus and Superfiltering Li et al. (2023a, 2024), involves a one-off instruction fine-tuning (One-off IFT) on the whole instruction dataset characterized by their diversity and high quality. This approach, while emphasizing dataset quality and diversity Wang et al. (2023b); Li et al. (2023a), neglects the inherent complexity of the instruction sets. Consequently, this one-off IFT fails to adequately equip models with the nuanced capability to comprehend and execute a wide array of instructions. Differs from the one-off IFT approach, our proposed phased IFT method initially segments the instruction dataset into multiple sub-datasets, arranged in a sequence from simple to complex in terms of instruction difficulty. Continuous uptraining is then executed across this sequence.
@@ 3 Method
The difficulty score histogram of the 52K Alpaca dataset is illustrated as shown in the left of Figure 2. Another interesting finding is that a majority of instructions in the 52K Alpaca are categorized as low difficulty, with approximately 63.5% scoring below 1.5. In contrast, within the Alpaca-clean dataset, instructions with a difficulty score lower than 1.5 account for 52.3%; in the LIMA dataset, the proportion of instructions with a difficulty score below 1.5 is merely 9.9%, while those scoring above 4.5 account for a significant 58.3%. Details of their comparison can be found in Appendix A.1. This highlights the feasibility of employing GPT-4 to assess the difficulty level of instructions.
@@ 4 Experiments
This paper seeks to validate the progressive alignment hypothesis, examining the relative efficacy of two instruction alignment training methodologies: One-off Instruction Fine-Tuning (One-off IFT) versus our proposed Phased Instruction Fine-Tuning (Phased IFT), through three distinct experiments. Firstly, Experiment 1 evidences Phased IFT’s superior performance over One-off IFT across two instruction data, establishing Phased IFT’s effectiveness. Secondly, Experiment 2 assesses the merit of stratifying instruction datasets by difficulty versus arbitrary segmentation, aiming to attribute Phased IFT’s success to difficulty-stratified instruction data categorization rather than to the phased training approach per se. Finally, Experiment 3, conducting a full permutation analysis of Phased IFT, seeks to demonstrate that the method’s success is specifically due to uptraining in a sequence of increasing difficulty, with alternative sequencing failing to produce comparable outcomes.
Datasets and Evaluation We conducted extensive experiments on the standard Alpaca 52K instruction data and its refined version Alpaca-cleaned, with the following three recent LLMs: Llama-2 7/13B and Mistral-7B respectively, evaluating them across 6 benchmark datasets, including Self-Instruct of size 252, WizardLM of size 218, Koala of size 180, Vicuna of size 80, OASST of size 188 Köpf et al. (2023), and Anthropic Bai et al. (2022) of size 129.
Following evaluation from Alpagasus Chen et al. (2024), we prompt GPT-4 to judge Win-Tie-Lose and compute a win rate. Under identical instruction and input, GPT-4 as an auto-grader Zheng et al. (2023); Chia et al. (2023) to score the quality of output from both One-off IFT and Phased IFT methods, on a scale of 1 to 10 (See prompt in Appendix A.2). For fairness, we score each instruction twice, i.e., placing Phased IFT’s output before/after the One-off IFT’s output. An instruction’s overall outcome is defined by:
@@
Stratifying Alpaca by Difficulty Score
Figure 4: Difficulty score histogram of Alpaca-cleaned data and its cumulative probability density.
@@
The instruction difficulty within the Alpaca and Alpaca-cleaned data are quantitatively assessed by GPT-4, assigning scores from 1 to 5, with higher score denoting increased complexity. Based on the density curve (e.g., Figure 2 for Alpaca and Figure 4 for Alpaca-cleaned), we adopt two threshold values, namely 1.5 and 3.5 for both Alpaca and Alpaca-cleaned, and segment them into three stages sub-datasets, respectively. We refer to as the Alpaca-3-stages and Alpaca-cleaned-3-stages.
@@
Additionally, to conduct ablation experiments, we also construct a randomly sampled 3-stages data, with each stage containing the same number of instructions as the difficulty-segmented subset. This is referred to as the Alpaca-rand-3-stages and Alpaca-cleaned-rand-3-stages data. Detailed statistics, as delineated in Table 1.
@@
Implementations We utilized Huggingface’s Trainer with DeepSpeed’s Zero3 for data-parallel, full-parameter fine-tuning in our training regimen. Consistently across experiments, a learning rate of 5e-6 and a duration of two training epochs were applied. We set a per-device batch size of 16 and a gradient accumulation steps of 1, a weight decay of 0.1, a warmup ratio of 0.1, and a cosine annealing learning rate scheduler.
4.1 Experiment 1: Comparison win rate of Phased IFT and One-off IFT
@@
Table 2: Comparisons of win rates between Phased IFT and One-off IFT. Phased IFT involves up-training on Alpaca-3-stages and Alpaca-cleaned-3-stages, while One-off IFT is fine-tuned solely on the original Alpaca and Alpaca-cleaned data. The changes in win rates are represented by percentage differences, where a "+" indicates an increase and "-" denotes a decrease in percentage.
@@
Base Models Self-Instruct WizardLM Koala Vicuna OASST Anthropic Data
Llama-2 7B +5.75 +10.32 +4.17 +11.25 +8.78 +12.40 Alpaca
Mistral-7B -0.20 -2.52 +6.67 +5.00 +3.19 +3.10
Llama 2-13B +4.37 +6.88 +10.56 +11.25 +10.11 +13.18
Llama 2-7B +3.77 +4.13 +4.17 +3.12 - - Alpaca-cleaned
Llama 2-13B +2.78 +3.21 +1.39 +6.88 +7.45 +9.30
Table 2 demonstrates that the Phased IFT method yields superior performance across a wide range of evaluation metrics compared to the One-off IFT method, with the sole exceptions being the Mistral-7B model’s lower win rates on the Self-Instruct and WizardLM benchmarks. Specifically, the Llama-2 7B model, when fine-tuned with the Phased IFT method on the Alpaca-3-stages instruction set, experienced a win rate increase ranging from a minimum of +4.17 absolute points (on Koala) to a maximum of +12.40 absolute points (on Anthropic). Similarly, the Llama-2 13B model’s win rate improved by at least +4.37 points (on Self-Instruct) to as much as +13.18 absolute points (on Anthropic). Notably, the Mistral-7B model saw a win rate increase of +6.67 absolute points on Koala and +3.10 points on Anthropic, but it exhibited a decrease in win rates of -0.20 and -2.52 absolute points on Self-Instruct and WizardLM, respectively.
On the Alpaca-cleaned instruction set, the win rate of the Llama 2-7B model fine-tuned via Phased IFT improved by at least +3.12 absolute points (on Vicuna) and up to +4.17 points (on Koala), while the Llama-2 13B model saw increases ranging from +1.39 absolute points (on Koala) to +9.30 points on Anthropic.
These results from Experiment 1 affirm the efficacy of the Phased IFT method for instruction-based fine-tuning of pre-trained LLMs, leading to notable performance enhancements on Llama-2 7B/13B and Mistral-7B models.
@@
4.2 Experiment 2: Comparison win rate of difficulty-stratified 3-stages and randomly sampled 3-stages
Table 3: Comparison of win rates between Phased IFT and One-off IFT methods is outlined in three rows. The first row (Row #1) assesses Phased IFT on Alpaca-rand-3-stages against One-off IFT using the original Alpaca 52K data. In the second row (Row #2), the comparison is made between Phased IFT on Alpaca-cleaned-rand-3-stages and One-off IFT on the original Alpaca-cleaned data. The third row (Row #3) evaluates the win rates of Phased IFT on Alpaca-3-stages in contrast to Phased IFT on Alpaca-rand-3-stages. Each entry corresponds to Table 2.
@@ 4.2 Experiment 2: Comparison win rate of difficulty-stratified 3-stages and randomly sampled 3-stages
Base Models Self-Instruct WizardLM Koala Vicuna OASST Anthropic Row
Llama-2 7B -1.19 -2.75 +0.56 +0.62 -0.66 +0.75 #1
Llama-2 13B -3.97 -3.67 -3.89 -3.75 -2.93 -3.49
Llama-2 7B -2.78 -2.52 -5.00 -1.88 - - #2
Llama-2 7B +9.13 +12.16 +13.61 +10.62 +5.85 +6.59 #3
Llama-2 13B +4.76 +7.57 +1.94 +10.00 +10.37 +10.85
Table 3 provides insight into the effectiveness of our Phased IFT method for fine-tuning pre-trained models Llama-2 7B and 13B on randomly sampled 3-stages instruction data. Notably, the Phased IFT method, in comparison to the One-off IFT approach, exhibited a decline in win rates across several test sets. Specifically, as listed in the Row #1 fine-tuned with Phased IFT on Alpaca-rand-3-stages, the win rate of the Llama-2 13B model decreased significantly compared to One-off IFT on the original Alpaca 52K data, with reductions ranging from -2.93 absolute points on Anthropic to -3.97 points on Self-Instruct.
The Row #2 of Table 3 which focuses on the application of Phased IFT for fine-tuning Llama-2 7B on Alpaca-cleaned-rand-3-stages, shows a pronounced decrease in win rate compared to the One-off IFT on the original Alpaca-cleaned data, with declines spanning from -1.88 absolute points on Vicuna to -5.00 absolute points on Koala.
Contrastingly, Row #3 of Table 3 demonstrates that applying Phased IFT for uptraining the Llama-2 7B and 13B models on Alpaca-3-stages, as opposed to on Alpaca-rand-3-stages, yielded significant improvements in win rate. Specifically, the Llama-2 7B model saw increases from +5.85 absolute points on OASST to +13.61 on Koala, while the Llama-2 13B model’s performance improved by +1.94 points on Koala and up to +10.85 points on Anthropic.
These results indicate that uptraining with randomly selected 3-stages instruction data does not inherently enhance performance and can lead to reductions across most test sets. This contrasts with experiment 1 where Phased IFT applied to difficulty-stratified 3-stages instruction sets, matched for instruction count and compared against the same baseline, yielded substantial performance gains. This highlights two key insights: firstly, the improvements observed are not solely attributable to the 3-stages uptraining process; secondly, categorizing instruction data by difficulty level provides a structured fine-tuning approach, enabling pre-trained models to progressively adapt to more complex tasks.
Table 4: Win rate of all permutations of uptraining of Llama-2 7B/13B on Alpaca-3-stages instruction data, compared to One-off IFT on the original Alpaca data. Each entry is the same as in Table 2.

### v2

@@ Abstract
Instruction Fine-Tuning enhances pre-trained language models from basic next-word prediction to complex instruction-following. However, existing One-off Instruction Fine-Tuning (One-off IFT) method, applied on a diverse instruction, may not effectively boost models’ adherence to instructions due to the simultaneous handling of varying instruction complexities. To improve this, Phased Instruction Fine-Tuning (Phased IFT) is proposed, based on the idea that learning to follow instructions is a gradual process. It assesses instruction difficulty using GPT-4, divides the instruction data into subsets of increasing difficulty, and uptrains the model sequentially on these subsets. Experiments with Llama-2 7B/13B/70B, Llama3 8/70B and Mistral-7B models using Alpaca data show that Phased IFT significantly outperforms One-off IFT, supporting the progressive alignment hypothesis and providing a simple and efficient way to enhance large language models. Codes and datasets from our experiments are freely available at https://github.com/xubuvd/PhasedSFT.
@@ 1 Introduction
Figure 1: In the context of increasing difficulty multi-stage sub-datasets, the trend of win rate growth for uptraining (Phased IFT) compared to One-off on the original dataset is observed. The gray horizontal line represents the performance baseline of One-off.
@@
To address the aforementioned issue, this paper introduces an instruction difficulty scoring mechanism. Specifically, the difficulty of an instruction is defined by the complexity of the instruction itself, the input, and the challenges involved in generating the output. We employ the strongest available models (such as GPT-4 OpenAI (2023)) as a teacher to score the difficulty of each instruction and input, as well as the challenge in generating the output, on a scale from 1 to 5, where higher scores indicate greater difficulty. Further, we plot probability density curve of the instruction difficulty as a heuristic guide Campello et al. (2020), coupled with expert judgment to select thresholds for difficulty scores. This approach enabled us to segment the instruction dataset into multi-stages sub-datasets, forming a sequence with increasing difficulty levels. Building upon this, we further propose the Progressive Alignment Hypothesis: Aligning a pre-trained model’s existing ability to predict the next word with the capability to generate content following human intent is a gradual learning process, progressively attaining alignment with human intent.
@@
We conducted extensive experiments using two leading open-source pre-trained models, Llama2 7/13/70B Touvron et al. (2023), Llama3 8/70B, and Mistral 7B Jiang et al. (2023), on two widely-used instruction datasets, Alpaca and its refined version AlpacaClean. To evaluate the training effectiveness, we utilized 6 benchmarks, with GPT-4 serving as the judge to measure the performance improvements of Phased IFT over One-off IFT in terms of win rate. Specially, Figure 1 illustrates the process wherein Phased IFT’s win rate progressively surpasses that of One-off IFT on the three-stage difficulty-increasing sub-datasets. It is evident that the three Llama2 models and the two Llama3 models exhibit a consistent trend of increasing win rates with the progression of uptraining on multi-stage sub-datasets, indicating the effectiveness of the progressive alignment hypothesis. Furthermore, we conducted three sets of experiments and an ablation study to confirm the effectiveness of Phased IFT. The main contributions of this paper include:
@@
  • We introduce the Progressive Alignment Hypothesis, which posits that aligning the generative capabilities of pre-trained models with human intent is a gradual process, rather than being achieved through a one-time fine-tuning on an instruction dataset. This hypothesis is supported by extensive experimental validation.
  • Extensive experiments have demonstrated that our proposed Phased IFT is more effective than One-off IFT. It has significantly outperformed One-off IFT across 6 benchmarks. Additionally, a series of ablation experiments have validated the effectiveness of the Progressive Alignment Hypothesis.
@@ 2 Related Work
Open-source Instruction Dataset The Alpaca dataset Taori et al. (2023) aims to improve instruction-following ability of LLMs, consisting of 52K instructions, generated using OpenAI’s text-davincic-003. It has improved the self-instruction framework Wang et al. (2023b), leading to a greater diversity in the Alpaca dataset compared to the Self-Instruction dataset. Alpaca-cleaned yahma (2023) is a quality-enhanced version of the original Alpaca dataset, which addresses hallucinations, empty outputs, and incorrect answers, thereby improving the overall quality of the dataset. In addition, LIMA Zhou et al. (2023) has also released a carefully curated dataset, including 1000 instructions, which encompasses 750 popular queries and responses from community forums such as Stack Exchange and wikiHow, as well as 250 meticulously crafted entries.
Instruction Fine-tuning Strategy The prevalent method of instruction fine-tuning, employed in models such as LIMA, Alpaca, Alpagasus and Superfiltering Li et al. (2023a, 2024), involves a one-off instruction fine-tuning (One-off IFT) on the whole instruction dataset characterized by their diversity and high quality. This approach, while emphasizing dataset quality and diversity Wang et al. (2023b), neglects the inherent complexity of the instruction sets. Consequently, this one-off IFT fails to adequately equip models with the nuanced capability to comprehend and execute a wide array of instructions. Differs from the one-off IFT approach, our proposed phased IFT method initially segments the instruction dataset into multiple sub-datasets, arranged in a sequence from simple to complex in terms of instruction difficulty. Continuous uptraining is then executed across this sequence.
@@ 3 Method
The difficulty score histogram of the 52K Alpaca dataset is illustrated as shown in the left of Figure 2. Another interesting finding is that a majority of instructions in the 52K Alpaca are categorized as low difficulty, with approximately 63.5% scoring below 1.5. In contrast, within the Alpaca-clean dataset, instructions with a difficulty score lower than 1.5 account for 52.3%; in the LIMA dataset, the proportion of instructions with a difficulty score below 1.5 is merely 9.9%, while those scoring above 4.5 account for a significant 58.3%. Details of their comparison can be found in Appendix A.1and A.2. This highlights the feasibility of employing GPT-4 to assess the difficulty level of instructions.
@@ 4 Experiments
This paper aims to validate the progressive alignment hypothesis by comparing the efficacy of two instruction training methodologies: One-off Instruction Fine-Tuning (One-off IFT) and our proposed Phased Instruction Fine-Tuning (Phased IFT), through three experiments and ablation studies. Experiment 1 demonstrates the superior performance of Phased IFT over One-off IFT across two instruction datasets, establishing its effectiveness. Experiment 2 evaluates the effectiveness of stratifying instruction datasets by difficulty versus arbitrary segmentation, aiming to attribute Phased IFT’s success to difficulty-stratified instruction data categorization rather than to the phased training approach. Experiment 3 conducts a full permutation analysis of Phased IFT to demonstrate that its success is specifically due to uptraining in a sequence of increasing difficulty, with alternative sequencing failing to produce comparable outcomes.
Datasets and Evaluation We conducted extensive experiments on the standard Alpaca 52K instruction data and its refined version, Alpaca-cleaned, using 6 recent LLMs: Llama-2 7/13/70B, Llama-3 8/70B and Mistral-7B. These models were evaluated across 6 benchmark datasets: Self-Instruction Wang et al. (2023b) of size 252, WizardLM Xu et al. (2024) 218, Koala Geng et al. (2023) 180, Vicuna Chiang et al. (2023) 80, OASST 188 Köpf et al. (2023), and Anthropic of size 129 Bai et al. (2022).
Following the evaluation framework established by Alpagasus Chen et al. (2024), we prompted GPT-4 to assess Win-Tie-Lose outcomes and compute win rates across the aforementioned 6 benchmarks. Subsequently, a weighted average win rate metric is calculated based on these 6 benchmark datasets, referred to as Avg. win rate. Using identical instruction and input, GPT-4 acts as an auto-grader Zheng et al. (2023); Chia et al. (2023) to score the quality of output from both One-off IFT and Phased IFT methods on a scale of 1 to 10 (see prompt in Appendix A.3). To ensure fairness, each instruction is scored twice, with Phased IFT’s output positioned before and after One-off IFT’s output. An instruction’s overall outcome is determined based on these evaluations.
@@
Stratifying Alpaca by Difficulty Score The instruction difficulty within the Alpaca and Alpaca-cleaned datasets is quantitatively assessed by GPT-4, assigning scores from 1 to 5, with higher scores indicating increased complexity. Using density curves (e.g., Figure 2 for Alpaca and Figure 5 for Alpaca-cleaned), we adopt threshold values of 1.5 and 3.5 for both datasets, segmenting them into three stages sub-datasets, referred to as Alpaca-3-stages and Alpaca-cleaned-3-stages.
@@
For ablation experiments, we construct randomly sampled 3-stages datasets, with each stage containing the same number of instructions as the difficulty-segmented subset. These are denoted as Alpaca-rand-3-stages and Alpaca-cleaned-rand-3-stages. Detailed statistics are provided in Table 1.
@@
Implementations In our training, we employed Huggingface’s Trainer coupled with DeepSpeed’s Zero3 for data-parallel, full-parameter fine-tuning. Consistently, a learning rate of 5e-6 and a 2 epoch duration were utilized across experiments. We set a per-device batch size of 4 for 70B model and 16 for other models, a gradient accumulation step of 1, with a weight decay of 0.1 and a warmup ratio of 0.1. Additionally, a cosine annealing learning rate scheduler was employed.
4.1 Experiment 1: compares the win rates of Phased IFT and One-off IFT
@@
Table 2: Comparisons of win rates of Phased IFT against One-off IFT. The changes in win rates are represented by percentage differences, where a "+" indicates an increase and "-" denotes a decrease in percentage.
@@
Base Models Self-Instruct WizardLM Koala Vicuna OASST Anthropic Avg. win-rate Data
Llama-2 7B +4.17 +6.42 +5.56 +6.88 +9.31 +14.34 +7.26 Alpaca
Mistral 7B -0.60 +5.96 +6.11 +16.25 +7.98 +12.02 +6.30
Llama-2 13B +1.98 +2.75 +7.78 +11.88 +11.44 +16.28 +7.35
Llama-2 70B -3.04 +2.57 +7.27 +13.46 +4.10 +0.00 +2.82
Llama-3 8B +3.05 +9.67 +9.52 +7.50 +6.32 +12.60 +7.64
Llama-3 70B +1.65 +10.10 +5.23 +5.84 +2.34 +7.87 +5.23
Llama-2 7B +4.76 +1.15 +4.44 +2.50 +5.85 +1.16 +3.53 Alpaca- cleaned
Llama-2 13B +4.56 +2.98 +5.83 +15.62 +6.91 +10.85 +6.49
Llama-2 70B +0.40 +6.42 +8.33 +12.50 +10.90 +14.73 +7.59
Llama-3 8B +5.75 +3.90 +1.15 +0.62 +4.79 +5.43 +3.97
Table 2 lists the win rate results of Phased IFT compared to the One-off IFT method, demonstrating that our proposed Phased IFT significantly outperforms the One-off IFT method across six benchmarks on two training datasets.
The first part of Table 2 compares the results of the two fine-tuning methods on Alpaca. Phased IFT was uptrained on Alpaca-3-stages, while One-off IFT was trained once on the original Alpaca. Using six base models, Phased IFT exceeded One-off IFT in the Avg. win-rate metric. Specifically, compared to the One-off IFT method trained once on the original Alpaca dataset, the Llama-2 7B using the Phased IFT fine-tuning method on Alpaca-3-stages showed a minimum win rate increase of +4.17 on Self-Instruct and a maximum of +14.34 on Anthropic, with an average win rate of +7.26. The Mistral 7B also achieved an Avg. win rate of +6.30; Llama 2 13B reached an average win rate of +7.35, with a win rate of +16.28 on Anthropic; Llama 3 8B and 70B both achieved an average win rate increase of over +5.0.
The second part of Table 2 compares the win rates of five base models using the two fine-tuning methods on Alpaca-cleaned. Specifically, the three Llama2 series models achieved average win rates of +3.53, +6.49, and +7.59, respectively, while Llama3 8B achieved an average win rate of +3.97. It can be observed that the average win rates of Llama2 7B/13B and Llama3 8B on Alpaca-cleaned are lower than their results on Alpaca, while Llama2 70B’s average win rate increased from +2.82 on Alpaca to +7.59 on Alpaca-cleaned. Due to time constraints, experiments on Llama3 70B have not yet been completed.
The results of Experiment 1 indicate that the Phased IFT fine-tuning method is indeed more effective than One-off IFT, achieving significantly higher win rates on most of the six benchmarks, with an overall average win rate significantly higher than that of One-off IFT.
@@
4.2 Experiment 2: Comparison of win rate between difficulty-stratified 3-stages and randomly sampled 3-stages
Table 3: Part #1 assesses win rates of Phased IFT on Alpaca-rand-3-stages against One-off IFT on the original Alpaca data. Part #2 evaluates win rates of Phased IFT on Alpaca-cleaned-rand-3-stages over One-off IFT on the original Alpaca-cleaned data. Part #3 measures win rates of Phased IFT on Alpaca-3-stages compared to Phased IFT on Alpaca-rand-3-stages.
@@
Base Models Self-Instruct WizardLM Koala Vicuna OASST Anthropic Avg. win-rate Part
Llama-2 7B -0.60 +2.98 -6.61 +3.12 -1.33 +1.94 -0.42 #1
Llama-2 13B -1.98 -2.29 -0.29 -6.88 -2.93 -0.39 -2.10
Llama-2 70B -10.32 -5.96 -6.90 -3.12 -11.97 -5.43 -7.96
Llama-3 8B -7.74 -1.38 -1.15 +5.00 +0.00 -2.33 -2.25
Llama-3 70B -10.00 -3.63 -4.44 -5.63 -6.89 -2.73 -5.92
Llama-2 7B -1.34 -5.07 -3.85 +8.12 -1.34 -2.76 -2.00 #2
Llama-2 13B +0.58 -0.69 +1.26 +1.88 +2.97 -0.39 0.84
Llama-3 8B +2.80 +0.69 -6.94 +2.50 +4.30 -6.59 -0.22
Llama-2 7B +2.38 +3.90 +11.78 +11.25 +2.39 +8.53 +5.74 #3
Llama-2 13B +6.94 +2.98 +6.32 +16.25 +11.70 +12.79 +8.29
Llama-2 70B +8.93 +5.50 +10.34 +19.38 +11.97 +8.14 +9.71
Llama-3 8B +2.18 +0.69 +13.79 +2.50 +8.78 +9.30 +5.95
Llama-3 70B +8.07 +14.10 +12.50 +17.12 +8.63 +5.96 +10.62
Table 4: Win rates of all permutations of Phased IFT compared to One-off IFT on the original Alpaca data.